OpenStudy (anonymous):

a hair salon gets a shipment of 84 bottles of hair conditioner to use and sell to customers.the two types of conditionders received are Type A, which is used for regular hair,and Type B which is used for dry Hair.Type A costst 6.50 per bottle and type b costs 8.25 per bottle.the hair salsons invoice for the conditioner is 588.how many of each type of conditioner are in the shipment? A)24 of type A,60 Type B B)7Type A,77Type B C)42Type A,42 Type B D)60Type A,24 Type B

OpenStudy (anonymous):

let x is the number of type A then no of type B becomes 84-x

OpenStudy (anonymous):

total cost = 6.5*x+8.25*(84-x)=588

OpenStudy (anonymous):

x=60

OpenStudy (anonymous):

do u understand

OpenStudy (anonymous):

so type A =x=60 and type B =84-x=84-60=24

OpenStudy (anonymous):

@sharee2012 do u understand...!!???????

OpenStudy (anonymous):

sorry im repsonding late but a little im going 2 attempt the pratice question 2 if i get it right ill let u knw if i get it wrong ill let u knw

OpenStudy (anonymous):

this is right 1000% no chance of it being wrong.....!!!!

OpenStudy (anonymous):

no i mean for me as far as understanding what u did

OpenStudy (anonymous):

ok if u dont understand u can always ask...!!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome...!!